HETAV PATEL

Grade 11,

The energy required to remove electron from sodium is 2.3 eV.Does sodium show photoelectric effect for orange light of wavelength 6800 A^o ?

The energy required to remove electron from sodium is 2.3 eV.Does sodium show photoelectric effect for orange light of wavelength 6800 A^o ?

Grade:12

1 Answers

Vikas TU
7 years ago
2.3 ev is thus the Work Functiongiven for Na.
from photoelectric eqn.
E = W + K.E.
Determine the E for wavelength 6800 eV.
Therfore,
E = hv = 1240/680 = 1.82 eV
which is hence less than Work Function 2.3 eV.
Therefore nO photoelctic effect would happen.